# Cupid: Pose-Grounded Generative 3D Reconstruction from a Single Image

### v1

@@ 2.3 Cascaded Flow Modeling
We employ a two-stage cascaded flow model to jointly sample a 3D object and its corresponding camera pose, i.e., $\mathbf{z}=\{(\mathbf{x}_{i},\mathbf{f}_{i}),(\mathbf{x}_{i},\mathbf{u}_{i})\}_{i=1}^{L}$. In the first stage, the flow model for occupancy and pose generation ($\mathbf{\mathcal{G}_{S}}$) generates (i) an occupancy cube that indicates active voxels and (ii) a UV cube that contains object-centric camera pose information (Section 2.2). In the second stage, conditioned on the predicted occupancy and camera pose, the pose-aligned geometry appearance generation model $\mathbf{\mathcal{G}_{L}}$ synthesizes DINO features $\mathbf{f}_{i}$ for each active voxel, yielding the final $\mathbf{z}$.
@@
In this stage, given a conditioning image $\mathbf{I}^{\textrm{cond}}$, we generate an occupancy cube and a UV cube with resolution $r$. The occupancy cube $\mathbf{G}_{o}\in\{0,1\}^{r\times r\times r\times 1}$ encodes binary values indicating active/nonactive voxels. The UV cube $\mathbf{G}_{uv}\in[0,1]^{r\times r\times r\times 2}$ stores normalized pixel coordinates $(u,v)$ for each voxel. Note that voxels sharing the same pixel coordinates are in the same ray [^3] We do not enforce occlusion as recent studies suggest that transformers can effectively model light transport (Jin et al., 2024; Zeng et al., 2025). Next, we train a 3D VAE to compress the UV cube into a low resolution feature grid $\mathbf{S}_{uv}\in\mathbb{R}^{s\times s\times s\times C}$ to improve computational efficiency. For fine-tuning TRELLIS, we concatenate the original feature grid $\mathbf{S}_{o}$ with $\mathbf{S}_{uv}$ and incorporate linear layers at both the input and output of the flow network $\mathbf{\mathcal{G}_{S}}$. Once the occupancy cube and the UV cube are generated, we can obtain $\{(\mathbf{x}_{i},\mathbf{u}_{i}(\bm{\theta}))\}_{i=1}^{L}$ by collecting active voxels and solving the camera pose with Equation 2.
@@
In this stage, we aim to generate the detailed 3D latent $\{\mathbf{f}_{i}\}_{i=1}^{L}$ only at active voxels. Our experiments show that the original $\mathbf{\mathcal{G}_{L}}$ from TRELLIS, which uses globally-attended image information, often suffers from color drift and loss of fine details (see Figure 8 in Appendix A.4). To address this issue, we leverage the calculated camera pose from the first stage to inject locally-attended pixel information in each voxel.
Specifically, we incorporate a single-view, pixel-aligned 3D latent derived from the input image and our generated camera pose to calculate the features located at the $i$-th voxel as follows:
@@
where each voxel’s pixel coordinate $\mathbf{u}_{i}$ is obtained by projecting the coordinates of the $i$-th 3D voxel center onto the image plane with the calculated camera pose, $\operatorname{BilinearInterp}$ denotes bilinear interpolation and $\operatorname{SlatEncoder}$ is the 3D VAE encoder. While DINO captures high-level semantics, it loses low-level cues needed for precise 3D geometry and appearance reconstruction. To compensate for it, we extract complementary low-level features $\mathbf{f}^{\mathrm{l}}$ from $\mathbf{I}^{\textrm{cond}}$ with a lightweight convolutional head and sample them at $\mathbf{u}_{i}$ via $\operatorname{BilinearInterp}$. Finally, for each voxel, we concatenate the current noisy voxel feature $\mathbf{f}_{i}^{t}$ at time step $t$ with the pixel-aligned features and fuse them into the flow transformer with a linear layer: $l_{t}=\mathrm{Linear}\big([\mathbf{f}_{i}^{t}\,\oplus\,\mathbf{f}^{\mathrm{h}}_{i}\,\oplus\,\mathbf{f}^{\mathrm{l}}_{i}]\big)$. This pose-aligned fusion significantly improves 3D geometry accuracy and appearance fidelity with respect to the input image.
3 Component-aligned Scene Reconstruction
Our pose-grounded generative framework seamlessly extends to scene-level compositional reconstruction by leveraging explicit 3D-to-2D correspondences for object-centric to view-centric transformations. We harness foundation models for object segmentation (Ravi et al., 2024), generate each 3D object, and compose them into a scene with a global depth prior (Wang et al., 2025d). Figure 3 shows one such result with our method. More results are given in Figure 9 of Appendix A.3.
Occlusion-aware 3D generation.
Images with multiple objects often contain mutual occlusions, which our synthetic training data do not model. To address this, we introduce random masks on the conditioning image during fine-tuning, inspired by Amodal3R (Wu et al., 2025). We adapt their masking strategy to our architecture; see Appendix A.2 for implementation. Then, given an input image with multiple objects and their corresponding object masks, we reconstruct each object independently with our occlusion-aware generator, yielding dense 3D–2D correspondences per object.
Multi-component composition.
Since absolute depth is not preserved across objects, direct 3D-2D alignment to the image is ill-posed. We address this using MoGe (Wang et al., 2025c) to predict a global pointmap, thereby reducing 3D–2D alignment to 3D–3D alignment. For the $k$-th object, we collect the matched pair $(\mathbf{x}_{i}^{k},\,\mathbf{u}_{i})$}, where $\mathbf{x}_{i}^{k}$ is the coordinates of a visible 3D point on the generated 3D object and $\mathbf{u}_{i}$ is the corresponding pixel coordinates within the mask $\mathbf{M}^{k}$. We query the MoGe pointmap to obtain $\mathbf{p}_{i}=P(\mathbf{u}_{i})$ in the camera frame, then estimate a per-object similarity transformation $\mathcal{S}_{k}=(s_{k},\mathbf{R}_{k},\mathbf{t}_{k})$ via the Umeyama method (Umeyama, 2002) on pairs $(\mathbf{x}_{i}^{k},\,\mathbf{p}_{i})$. Applying $\mathcal{S}_{k}$ places each object in the shared camera frame, yielding a component-aligned scene reconstruction. This composition can also extend to multi-view input using VGGT (Wang et al., 2025a), allowing for more flexible input and transforming partial geometry into holistic geometry.
4 Experiments
We evaluate the most relevant baselines on three tasks: monocular geometry prediction, input-view consistency, and single-image-to-3D reconstruction quality in Section 4.1. We do not include per-scene optimization methods, focusing on learning-based systems. We also ablate different pose-conditioning designs in Section 4.2. Implementation details are in Appendix A.1. For 3D generation methods that lack camera poses, we provide qualitative comparisons in the Appendix A.4.
4.1 Evaluation
Baselines.
We compare against three complementary families for single image 3D reconstruction: point-map regression, view-centric 3D reconstruction, and 3D generation with post-hoc pose alignment. Each probes a distinct capability: visibility-limited lifting, view-centric full 3D recovery, and decoupled canonical 3D object generation and pose estimation.
Point-map regression. VGGT (Wang et al., 2025a) and MoGe (Wang et al., 2025c) simultaneously predict per-pixel 3D points and view-centric camera pose for 2D-to-3D reconstruction. These methods reconstruct only the geometry visible in the input view, often lacking robust priors for occluded regions. View-centric 3D reconstruction. LRM (Hong et al., 2023) and LaRa (Chen et al., 2024) generate view-centric 3D models, eliminating the need to estimate camera parameters at test time. LRM directly regresses 3D objects in view space, while LaRa enhances the input with novel views from Zero123++ (Shi et al., 2023) before reconstruction. We adopt an open-source implementation of LRM (He & Wang, 2023) for comparisons. 3D generation with post-hoc alignment. OnePoseGen (Geng et al., 2025) integrates a learned 3D generator (Ye et al., 2025), which produces canonical 3D models, with a robust object pose estimator (Wen et al., 2024) to align the 3D model with the input view. This approach evaluates whether decoupling 3D generation from pose estimation and applying post-hoc alignment can achieve reprojection consistency.
Evaluation.
We evaluate three aspects: (i) monocular geometry accuracy (following (Wang et al., 2025c)), (ii) input-view consistency to assess reprojection alignment and appearance fidelity, and (iii) novel-view consistency to assess full 3D fidelity. Experiments are conducted on Toys4K (Stojanov et al., 2021) and GSO (Downs et al., 2022), containing approximately 3K synthetic and 1K real-world objects, respectively. For novel views, we render four uniform views rotating around the vertical axis. Note that tasks (i) and (ii) are particularly challenging for generative approaches, as they require producing an object-centric camera pose for view-centric evaluation.
Monocular geometry accuracy.
Monocular geometry evaluates visible geometry in the input image. We construct ground-truth point clouds by unprojecting depth maps with camera parameters. We report Mask IoU, Chamfer Distance (CD), and F-score to assess point-cloud quality, providing both mean and median statistics to mitigate outlier sensitivity in generative methods. Since VGGT outputs point maps without object masks, we use ground-truth masks to avoid background contamination. For full 3D methods, we render and unproject the depth map to get the visible point cloud, and use rendered alphas as the object mask. Predicted point clouds are aligned to ground truth using the scale–shift alignment from MoGe (Wang et al., 2025c). Results are in Table 1.
@@ 4.1 Evaluation
Table 1: Monocular geometry accuracy. Cupid outperforms all 3D reconstruction and generation baselines and matches point-map regression methods that predict only partial geometry. VGGT uses a ground-truth object mask, which may overestimate accuracy.
@@
Figure 4: Comparison of generative reconstruction from a single image. Our method produces the highest-fidelity geometry and appearance; LRM hallucinates incorrect details, LaRa is overly blurry due to 2D diffusion inconsistencies, and OnePoseGen frequently fails to register pose reliably.
Our approach significantly outperforms full 3D reconstruction methods across all metrics. Notably, it reduces the average CD from LRM by approximately 50% on the GSO dataset, with even greater improvements in median statistics. These results highlight the superiority of our canonical object-centric reconstruction compared to the view-centric approaches of LRM and LaRa. OnePoseGen falls short on Toys4K and GSO because these datasets contain geometric symmetries and textureless regions that cause its registration modules to fail. Our method is also highly competitive with VGGT and MoGe across all metrics, though these methods only produce partial geometries.
Table 2: Input-view consistency. Cupid achieves superior input view consistency, producing accurate pose, geometry and appearance alignment.
@@
Input view consistency.
We evaluate the input view consistency to measure the alignment between the 3D reconstruction and the input image. We quantify the difference between the re-rendered image and the input image, reporting PSNR, SSIM, and LPIPS in Table 2. We also provide qualitative comparisons in Figure 4. Our method achieves substantial PSNR improvements over existing reconstruction-based methods: 36% and 44% over LaRa on Toys4K and GSO respectively, and 13% and 11% over LRM on the same datasets. We find that LaRa’s texture quality is affected by inconsistencies in the generated novel views, while LRM struggles to predict detailed textures. OnePoseGen can produce plausible 3D objects but suffers severely from texture misalignment (e.g., color shifting) and imperfect pose estimation. Our method achieves strong pose accuracy while retaining high-quality texture details, as evidenced by the zoomed-in views in Figure 4.
Full 3D evaluation.
Evaluating single-image-to-3D is challenging because many distinct 3D shapes can explain a single view. We therefore emphasize (1) full-3D qualitative assessments (Figure 4) and (2) quantitative semantic evaluations using CLIP (Radford et al., 2021) on Toys4K (Table 3). Our method consistently outperforms all baselines. Moreover, it reconstructs high-quality 3D objects and scenes from images in the wild (Figure 1), which is not achieved by prior methods.
4.2 Ablation Studies
Table 3: Comparison on CLIP image scores of novel views.
Method ViT-B/16 ViT-L/14
OnePoseGen 0.7933 0.7193
LaRa 0.8334 0.7682
OpenLRM 0.8939 0.8410
TRELLIS 0.9465 0.9210
Ours 0.9501 0.9291
@@ 4.2 Ablation Studies
We compare five pose-aligned conditioning variants on Toys4K: (a) the TRELLIS baseline; (b) adding DINOv2 positional embeddings to SLat latents; (c) concatenating DINOv2 feature volumes with view-conditioned voxel latents from the frozen SLat encoder; (d) as in (c) but zeroing latents of voxels occluded in the input view; and (e) concatenating additional visual features from a convolution layer applied to the input image before DINOv2 downsampling. From Table 4, method (b) and (c) validate the benefit of pose-aligned conditioning in generating geometry and appearance; (d) indicates that performance is equal whether occlusion is present or absent; our full method (e) provides the best visual quality by supplementing DINOv2 features with missing low-level cues, yielding better color and detail alignment, as shown in Figure 5. These gains persist with sampled geometry and pose, indicating our pose-aligned feature injection is robust to small variations.
5 Related Work
3D reconstruction from many images.
Complete 3D object reconstruction traditionally requires multiple views and Structure-from-Motion (Schönberger & Frahm, 2016; Schönberger et al., 2016). DUSt3R (Wang et al., 2024) predicts pixel-aligned point maps for image pairs, enabling efficient recovery of view-centric camera poses and partial geometry; subsequent works improve efficiency and flexibility (Wang et al., 2025a; d; b; Zhang et al., 2025; Yang et al., 2025a; Karaev et al., 2024; Yang et al., 2024; Liu et al., 2024d; Cabon et al., 2025). CUT3R processes images recurrently (Wang et al., 2025b); VGGT jointly predicts depth, point maps, and poses (Wang et al., 2025a); and MoGe introduces affine-invariant point maps for monocular geometry (Wang et al., 2025c). Unlike 2D-to-3D regression, our method generates 2D coordinates from 3D volumes, allowing for complete recovery of 3D objects and object-centric camera poses.
3D reconstruction from one or a few images.
High-quality multiview data are often hard to obtain in practice, motivating single- or sparse-view 3D reconstruction with large models (Xu et al., 2024a; Wei et al., 2024; Liu et al., 2024b; Tochilkin et al., 2024; Tang et al., 2024; Yinghao et al., 2024; Xu et al., 2024c). LRM (Hong et al., 2023) shows that complete 3D can be regressed from a single image under a fixed camera. PF-LRM (Wang et al., 2023) enables pose-free sparse-view reconstruction via differentiable PnP (Chen et al., 2022). Subsequent work improves single-view quality by augmenting novel views from 2D generative priors and using more efficient representations (Xu et al., 2024b; Yinghao et al., 2024; Liu et al., 2023a) or architectures (Chen et al., 2024). However, all methods are view-centric, not object-centric. Furthermore, these methods cannot generate diverse and plausible 3D objects that accurately explain the input image.
3D reconstruction from generative priors.
Dreamfusion (Poole et al., 2022) pioneer 3D generation from 2D diffusion models (Rombach et al., 2021; Saharia et al., 2022). Following work Zero-1-to-3 (Liu et al., 2023b) fine-tune image diffusion models for novel view synthesis, enabling single image-to-3D generation and inspiring follow-ups (Shi et al., 2023; Watson et al., 2022; Li et al., 2024; Liu et al., 2024c; Gao et al., 2024; Höllein et al., 2024; Wu et al., 2024). Recent methods (Zhou et al., 2025; Ren et al., 2025; Song et al., 2025) utilize video diffusion priors to achieve greater view consistency. Yet, all these methods, which leverage 2D priors, require multi-view reconstruction to obtain the 3D object, sharing similar limitations of many images to 3D reconstruction. The very recent attempts (Huang et al., 2025; Yao et al., 2025; Dong et al., 2025) incorporate 3D generative priors (Xiang et al., 2025; Zhang et al., 2024b) for scene reconstruction, but relies on costly ad-hoc optimization (Yao et al., 2025) or extra scene datasets (Huang et al., 2025).
@@ 5 Related Work
Figure 6: Multi-view conditioning. When multiple input views are available, we fuse the shared object latent across flow paths (similar to MultiDiffusion (Bar-Tal et al., 2023)), enabling camera, geometry and texture refinement across all images. Top: inputs; Middle: reconstructed 3D object and camera poses; Bottom: rendered images and geometry.

### v2

@@ 3.3 Cascaded Flow Modeling
We employ a two-stage cascaded flow model to jointly sample a 3D object and its corresponding camera pose, i.e., $\mathbf{z}=\{(\mathbf{x}_{i},\mathbf{f}_{i}),(\mathbf{x}_{i},\mathbf{u}_{i})\}_{i=1}^{L}$. In the first stage, the flow model for occupancy and pose generation ($\mathbf{\mathcal{G}_{S}}$) generates (i) an occupancy cube that indicates active voxels and (ii) a UV cube that encodes object-centric camera pose (Sec. 3.2). In the second stage, conditioned on the predicted occupancy and camera pose, the pose-aligned geometry and appearance generation model $\mathbf{\mathcal{G}_{L}}$ synthesizes DINO features $\mathbf{f}_{i}$ for each active voxel, yielding the final $\mathbf{z}$.
@@
In this stage, given a conditioning image $\mathbf{I}^{\textrm{cond}}$, we generate an occupancy cube and a UV cube with resolution $r$. The occupancy cube $\mathbf{G}_{o}\in\{0,1\}^{r\times r\times r\times 1}$ encodes binary values indicating active/inactive voxels. The UV cube $\mathbf{G}_{uv}\in[0,1]^{r\times r\times r\times 2}$ stores normalized pixel coordinates $(u,v)$ for each voxel. Voxels sharing the same pixel coordinates lie on the same rayWe do not model occlusion as recent studies suggest that transformers can effectively model light transport [jin2024lvsm, zeng2025renderformer].. Next, we train a 3D VAE to compress the UV cube into a low-resolution feature grid $\mathbf{S}_{uv}\in\mathbb{R}^{s\times s\times s\times C}$, improving efficiency with near-lossless pose accuracy (mean RRE / RTE << 0.5°). For fine-tuning TRELLIS, we concatenate the original feature grid $\mathbf{S}_{o}$ with $\mathbf{S}_{uv}$ and incorporate linear layers at both the input and output of the flow network $\mathbf{\mathcal{G}_{S}}$. Once the occupancy and UV cubes are generated, we obtain $\{(\mathbf{x}_{i},\mathbf{u}_{i}(\bm{\theta}))\}_{i=1}^{L}$ by collecting active voxels and solving the camera pose with Eq. 2.
@@
In this stage, we generate the detailed 3D latent $\{\mathbf{f}_{i}\}_{i=1}^{L}$ only at active voxels. Our experiments show that the original $\mathbf{\mathcal{G}_{L}}$ from TRELLIS, which uses globally-attended image information, often suffers from color drift and loss of fine details (see Fig. 9 in App. A.4). To address this, we leverage the calculated camera pose from the first stage to inject locally attended pixel information in each voxel.
Specifically, we leverage the estimated camera pose to compute the features located at the $i$-th voxel as follows:
@@
where each voxel’s pixel coordinate $\mathbf{u}_{i}$ is obtained by projecting the coordinates of the $i$-th 3D voxel center onto the image plane with the calculated camera pose, $\operatorname{BilinearInterp}$ denotes bilinear interpolation and $\operatorname{SlatEncoder}$ is the 3D VAE encoder. While DINO captures high-level semantics, it loses low-level cues needed for precise 3D geometry and appearance reconstruction. To compensate, we extract complementary low-level features $\mathbf{f}^{\mathrm{l}}$ from $\mathbf{I}^{\textrm{cond}}$ with a lightweight convolutional head and sample them at $\mathbf{u}_{i}$ via $\operatorname{BilinearInterp}$. Finally, for each voxel, we concatenate the current noisy voxel feature $\mathbf{f}_{i}^{t}$ at time step $t$ with the pixel-aligned features and fuse them into the flow transformer with a linear layer: $l_{t}=\mathrm{Linear}\big([\mathbf{f}_{i}^{t}\,\oplus\,\mathbf{f}^{\mathrm{h}}_{i}\,\oplus\,\mathbf{f}^{\mathrm{l}}_{i}]\big)$. This pose-aligned fusion significantly improves 3D geometry accuracy and appearance fidelity with respect to the input image.
@@
Table 1: Monocular geometry accuracy. Cupid outperforms all 3D reconstruction and generation baselines and matches point-map regression methods that predict only partial geometry. Note that VGGT uses a ground-truth object mask, which may overestimate accuracy.
@@
Figure 5: Qualitative comparison on input view consistency. We render the input view using its generated camera pose. For view centric methods (LRM, LaRa), we use ground-truth intrinsic for rendering as they do not model intrinsic. Our method produces the highest-fidelity geometry and appearance; LRM hallucinates incorrect details, LaRa is overly blurry due to 2D diffusion inconsistencies, and 3D generation method OnePoseGen frequently fails to register pose reliably.
4 Component-aligned Scene Reconstruction
By explicitly modeling the spatial relationship between the generated object and the input camera pose, our approach seamlessly integrates powerful generative priors into 3D reconstruction. Cupid naturally extends to scene-level compositional reconstruction by using explicit 3D-to-2D correspondences for spatial alignment. We first harness foundation models for object segmentation [ravi2024sam], generate each 3D object, and compose them into a scene with a global depth prior [wang2025moge2]. Fig. 4 shows one such result with our method. More results are given in Fig. 10 of App. A.3.
Occlusion-aware 3D generation.
Images with multiple objects often contain mutual occlusions, which our synthetic training data do not model. To address this, we introduce random masks on the conditioning image during fine-tuning, inspired by Amodal3R [wu2025amodal3r]. We adapt their masking strategy to our architecture; see App. A.2 for implementation. Then, given an input image with multiple objects and their corresponding object masks, we reconstruct each object independently with our occlusion-aware generator, yielding dense 3D–2D correspondences per object.
Multi-component composition.
Since absolute depth is not preserved across objects, direct 3D-2D alignment to the image is ill-posed. We address this using MoGe [wang2025moge] to predict a global pointmap, thereby reducing 3D–2D alignment to 3D–3D alignment. For the $k$-th object, we collect the matched pair $(\mathbf{x}_{i}^{k},\,\mathbf{u}_{i})$}, where $\mathbf{x}_{i}^{k}$ is the coordinates of a visible 3D point on the generated 3D object and $\mathbf{u}_{i}$ is the corresponding pixel coordinates within the mask $\mathbf{M}^{k}$. We query the MoGe pointmap to obtain $\mathbf{p}_{i}=P(\mathbf{u}_{i})$ in the camera frame, then estimate a per-object similarity transformation $\mathcal{S}_{k}=(s_{k},\mathbf{R}_{k},\mathbf{t}_{k})$ via the Umeyama method [umeyama2002least] on pairs $(\mathbf{x}_{i}^{k},\,\mathbf{p}_{i})$. Applying $\mathcal{S}_{k}$ places each object in the shared camera frame, yielding a component-aligned scene reconstruction. This composition can also extend to multi-view input using VGGT [wang2025vggt], allowing for more flexible input and transforming partial geometry into holistic geometry.
5 Experiments
We evaluate key baselines on three tasks: monocular geometry prediction, input-view consistency, and single-image-to-3D reconstruction in Sec. 5.1. We exclude per-scene optimization methods and focus on learning-based systems. We ablate pose-conditioning in Sec. 5.2, with implementation details and pose encoding accuracy in App. A.1. We do not extensively compare with state-of-the-art 3D generators, as improving their aesthetic or geometric quality is orthogonal to our goal. Qualitative comparisons with our baseline [xiang2025structured] for input-view consistency appear in Sec. A.4.
5.1 Evaluation
Baselines.
We compare against three complementary families for single image 3D reconstruction: point-map regression, view-centric 3D reconstruction, and 3D generation with post-hoc pose alignment.
Table 2: Input-view consistency. Cupid achieves superior input view consistency, producing accurate appearance alignment.
@@ 5.1 Evaluation
Table 3: Comparison on full 3D quality. We report CLIP image scores of novel views following [gao2024cat3d].
OnePoseGen LaRa OpenLRM TRELLIS Ours
ViT-B/16 0.7933 0.8334 0.8939 0.9465 0.9501
ViT-L/14 0.7193 0.7682 0.8410 0.9210 0.9291
Point-map regression. VGGT [wang2025vggt] and MoGe [wang2025moge] simultaneously predict per-pixel 3D points and view-centric camera pose for 2D-to-3D reconstruction. These methods reconstruct only the geometry visible in the input view, often lacking robust priors for occluded regions. View-centric 3D reconstruction. LRM [hong2023lrm] and LaRa [LaRa] generate view-centric 3D models, eliminating the need to estimate camera parameters at test time. LRM directly regresses 3D objects in view space, while LaRa enhances the input with novel views from Zero123++ [shi2023zero123++] before reconstruction. We adopt an open-source implementation of LRM [openlrm] for comparisons. 3D generation with post-hoc alignment. OnePoseGen [geng2025one] integrates Higen3D [ye2025hi3dgen] and TRELLIS [xiang2025structured], which produces canonical 3D models, with FoundationPose [wen2024foundationpose] to align the 3D model with the input view. This approach evaluates whether decoupling 3D generation from pose estimation and applying post-hoc alignment can achieve reprojection consistency.
Evaluation.
We evaluate three aspects: (i) monocular geometry accuracy, (ii) input-view consistency to assess reprojection alignment and appearance fidelity, and (iii) novel-view consistency to assess full 3D fidelity. Experiments are conducted on Toys4K [stojanov2021using] and GSO [downs2022google], containing approximately 3K synthetic and 1K real-world objects, respectively. For novel views, we render four uniform views rotating around the vertical axis. Tasks (i) and (ii) are particularly challenging for generative approaches, as they require producing accurate an object-centric camera pose that aligns with the input view.
Monocular geometry accuracy.
Monocular geometry metrics evaluate pose alignment and visible geometry fidelity. Unlike multi-view approaches that report pose accuracy relative to the first frame, single-view generation cannot report relative pose metrics. Following MoGe [wang2025moge], we use monocular geometry alignment as a proxy for pose accuracy. We construct ground-truth (GT) point clouds by unprojecting depth maps with camera parameters and report Mask IoU, Chamfer Distance (CD), and F-score, providing both mean and median statistics to mitigate outlier sensitivity. Since VGGT outputs point maps without masks, we use GT masks to avoid background contamination. For full 3D methods, we render and unproject the depth map to obtain the visible point cloud, using rendered alphas as the mask. Predicted point clouds are aligned to GT using scale–shift alignment from MoGe [wang2025moge]. Results are in Tab. 1.
@@
Figure 7: Multi-view conditioning. Our decoupled joint modeling naturally supports multi-view conditioning. When multiple input views are available, we fuse the shared view-agnostic object latent across flow paths (similar to MultiDiffusion [bar2023multidiffusion]), enabling object and cameras refinement across all views. Top: inputs; Middle: reconstructed 3D object and camera poses; Bottom: rendered images and geometry.
Our approach demonstrates clear superiority over view-centric reconstruction methods such as LRM and LaRa. On the GSO dataset, it reduces the average Chamfer Distance (CD) by 50% relative to LRM. This improvement stems not only from accurate intrinsic modeling via generative priors but also from our precise extrinsic camera parameter prediction. By contrast, although OnePoseGen excels in intrinsic modeling, it falls short on Toys4K and GSO, as these datasets contain geometric symmetries and textureless regions that cause its registration modules to fail during extrinsic pose estimation. Crucially, our model is also highly competitive with methods that produce only partial geometries (e.g., VGGT, MoGe). This suggests that complex, multi-stage pipelines which use these partial methods as a first step [yao2025castcomponentaligned3dscene] can be streamlined. Our approach achieves comparable monocular accuracy in a single step, eliminating the need for such intermediate representations.
Input view consistency.
We evaluate input view consistency to measure alignment between the 3D reconstruction and the input image. We quantify the difference between the re-rendered and input images, reporting PSNR, SSIM, and LPIPS in Tab. 2, with qualitative comparisons in Fig. 5. Our method achieves substantial fidelity improvements over existing reconstruction-based methods. We find LaRa’s texture quality is affected by inconsistencies in the generated novel views, while LRM struggles to predict detailed textures. OnePoseGen can produce plausible 3D objects but suffers severely from texture misalignment (e.g., color shifting), often failing on pose estimation. Our method achieves strong pose accuracy while retaining high-quality texture details, as evidenced by the zoomed-in views in Fig. 5.
Full 3D evaluation.
Evaluating single-image-to-3D is challenging because many distinct 3D shapes can explain a single view. We therefore emphasize (1) qualitative assessments (Fig. 5) and (2) quantitative semantic alignment using CLIP similarity [radford2021learning] on Toys4K (Tab. 3). Our method consistently outperforms all baselines. Moreover, it reconstructs high-fidelity 3D objects and scenes from in-the-wild images (Fig. 2), which prior methods do not achieve.
5.2 Ablation Studies
We compare pose-aligned conditioning (PAC) variants in the second-stage refinement on Toys4K: (a) the TRELLIS baseline (without PAC); (b) adding DINOv2 positional embeddings to SLat latents; (c) concatenating DINOv2 feature volumes with view-conditioned voxel latents from the frozen SLat encoder; (d) as in (c) but zeroing latents of occluded voxels in the input view; and (e) concatenating additional visual features from a convolutional layer applied to the input image. We report results using both GT coarse geometry and camera pose, and samples from the first stage.
As shown in Tab. 4, when provided with ground-truth coarse geometry and camera pose, methods (b) and (c) validate the benefit of pose-aligned conditioning in generating geometry and appearance; (d) indicates that performance remains consistent whether occlusion is present or absent; our full method (e) provides the best visual quality by supplementing DINOv2 features with missing low-level cues, yielding better color and detail alignment, as shown in Fig. 6. Crucially, these advantages are largely preserved when using sampled geometry and camera poses from the first stage (30.05 PSNR for (e) vs. 27.47 for baseline (a)). This confirms that our pose-aligned conditioning is highly robust to small perturbations from stochastic generation.